if log10x=-2, what is the value of x?
Answer by dabanfield(803)

if log10x=-2, what is the value of x?
If log 10x = -2 then by definiiton of a log:
10^(-2) = 10x
1/10^2 = 10x
1/100 = 10x
x = (1/100)/10
x = 1/1000